OpenStudy (brittany16):

Solve the equation. M over -6+6=41?

OpenStudy (anonymous):

Wolfram. Alpha.

OpenStudy (anonymous):

-210, unless it's \[m \div(-6+6)=41\]

OpenStudy (anonymous):

Don't be rude. We have answered about 10 or so questions exactly like this. Why don't you type out what you have so far and we can try to help you?

OpenStudy (anonymous):

Alright, sorry. I guess it's helpful to post stuff here, because WA only gives you the final answer without showing you the steps.

OpenStudy (brittany16):

dude you havent asnswer pellet so stop pleayin your lil games owletgot

OpenStudy (anonymous):

Uh huh... so what you're saying is you don't have anything to show us? So how can we help you figure out how to solve the problem if we don't know what kind of trouble you're having?

OpenStudy (anonymous):

How does one answer a pellet? And what is a owletgot? lolwut?

OpenStudy (shadowfiend):

Watch your language brittany. The fact that we have filters doesn't mean you can just say whatever you want.

OpenStudy (shadowfiend):

DS -- if you don't want to help, don't help. No one is forcing you to answer all of brittany16's questions.

OpenStudy (shadowfiend):

And last but not least, please don't go down the list and report 8 million posts, even if they're from the same person. We'll get the abuse report once, we don't need 80 emails to come check it out ;)

OpenStudy (anonymous):

I want to help, but I don't want to simply solve her questions. I want to teach her how to fish; I don't want to keep catching her a fish over and over. So I asked her to put what she had so far

OpenStudy (shadowfiend):

With all that said... Brittany, I think the helpers on the site would appreciate it if you at least worked with them to solve the problems, rather than just taking answers.

OpenStudy (shadowfiend):

It'll help you more anyway when you end up in a situation where you can't use OpenStudy :)

jhonyy9 (jhonyy9):

(M/-6)+6=41 M/-6=35 M=35*(-6) M= -210